A circular garden, 20 ft in diameter, has a path around it that is 3 ft wide. Find the area of the path? How much fence is needed if the garden is fenced (not including the path)?
Answer by TimothyLamb(4379)

---
area of garden and path:
A = pi*13*13
---
area of garden only:
a = pi*10*10
---
area of path:
A - a = pi*13*13 - pi*10*10
A - a = 216.77 sq.ft
---
garden fencing needed:
c = 2*pi*10
c = 62.83 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php